# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v2

@@ Abstract
Recent advancements in visualizing deep neural networks provide insights into their structures and mesh extraction from Continuous Piecewise Affine (CPWA) functions. Meanwhile, developments in neural surface representation learning incorporate non-linear positional encoding, addressing issues like spectral bias; however, this poses challenges in applying mesh extraction techniques based on CPWA functions. Focusing on trilinear interpolating methods as positional encoding, we present theoretical insights and an analytical mesh extraction, showing the transformation of hypersurfaces to flat planes within the trilinear region under the eikonal constraint. Moreover, we introduce a method for approximating intersecting points among three hypersurfaces contributing to broader applications. We empirically validate correctness and parsimony through chamfer distance and efficiency, and angular distance, while examining the correlation between the eikonal loss and the planarity of the hypersurfaces.
@@ 1 Introduction
Recent advancements in visualizing deep neural networks [1, 2, 3, 4] significantly contribute to understanding their intricate structures. This progress provides valuable insights into the expressivity, robustness, training methodologies, and distinctive geometry of neural networks. By leveraging the inherent piecewise linearity in the Continuous Piecewise Affine (CPWA) functions, e.g., ReLU neural networks, each region of the input space is represented as a convex polyhedral. The assembly of these sets constructs a polyhedral complex that delineates the decision boundaries of neural networks [5].
@@
    This provides a theoretical exposition of the eikonal constraint, revealing that within the trilinear region, the hypersurface transforms into a plane.
@@ 2 Related work
A conventional black-box approach is to assess a given function at sampled points to derive the complex of the function, commonly through marching cubes [9] or marching tetrahedra [16, 10]. Despite the inclusion of mesh optimization techniques [17] in modern mesh packages, this exhaustive method is still inclined to generate redundant complexities due to the discretization, needlessly burdening the computational load on a learning pipeline. Aiming for exact extraction, initial efforts focused on identifying linear regions [18] and employing Analytical Marching [2] to exactly reconstruct the zero-level isosurface. Contemporary approaches involve region subdivision, wherein the regions of the complex are progressively subdivided from neuron to neuron and layer to layer, allowing for the efficient calculation of the exponentially increasing linear regions [19, 20, 21, 4]. Nevertheless, many previous studies have explored CPWA functions consisting of fully connected layers and ReLU activation, primarily owing to the mathematical simplicity of hyperplanes defining linear regions. However, this restricts their applications in emerging domains of deep implicit surface learning that diverge from CPWA functions.
@@
The spectral bias of a multi-layer perceptron (MLP) hinders effectively learning high frequencies, both theoretically and empirically [14]. To address this limitation, Fourier feature mapping, employing trigonometric projections similar to those used in the Transformer architecture [22], proves successful in representing complex 3D objects and scenes. Furthermore, to address its slow convergence, positional encoding methods and enhance rendering quality, novel positional encoding methods using trilinear interpolations, e.g., HashGrid [12], TensoRF [13] are introduced. Both techniques generate positional features through trilinear interpolation among the eight nearest corner features derived from a pre-defined 3D grid. These features are obtained by hashing from multi-resolution hash tables or factorized representations of a feature space, respectively. However, these approaches cause the function represented by neural networks to deviate from CPWA functions, as trilinear interpolation is not an affine transformation. Consequently, the complex regions are divided by intricate hypersurfaces.
@@
An eikonal  [^1] The eikonal is derived from the Greek, meaning image or icon. equation is a non-linear first-order partial differential equation governing the propagation of wavefronts. Notably, this equation finds application in the regularization of a signed distance function (SDF) in the context of neural surface modeling [23, 6, 8]. If $\Omega$ represents a subset of the space ${\mathcal{X}}$ equipped with the metric $d$, the SDF $f({\bm{x}})$ is defined as follows:
@@
where $\partial\Omega$ denotes the boundary of $\Omega$. The metric $d$ is then expressed as:
@@ 3.1 Tropical geometry
Tropical  [^2] The tropical is named in honor of Hungarian-born Brazilian computer scientist Imre Simon [24]. geometry [25, 26] describes polynomials and their geometric characteristics. As a skeletonized version of algebraic geometry, tropical geometry represents piecewise linear meshes using tropical semiring and tropical polynomial (ref., Section A.1). The set of points where a tropical polynomial $f$ is non-differentiable, due to tropical sum $\oplus:=\max(\cdot)$, is called tropical hypersurface:
@@
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$ and tropical product :=$\odot$:=+⊙ := +.
@@
Without loss of generality, $L$-layer neural networks with ReLU activation are a composition of functions defined as:
@@
Although the neural networks $\nu$ is generally non-convex, it is known that the difference of two tropical signomials can represent $\nu$ [27] (ref., Section A.3). Given this observation, we proceed to examine the following proposition:
Proposition 3.3.
(Decision boundary of neural networks) The decision boundary for $L$-layer neural networks with ReLU activation $\nu^{(L)}({\bm{x}})=0$ is where two tropical monomials, or two arguments of $\max$, equal (Definition 3.1), recursively, as follows:
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)\big{\}}$
where the subscription $j$ denotes the $j$-th element, assuming the hidden size of neural networks is $H$, while the final $L$-th layer has a single output for rather discussions as an SDF. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
Given that piecewise hyperplanes corresponding to the preactivations, $\nu_{j}^{(i)}$ are the candidates shaping the decision boundary, it enables for-loop iterations over the neurons across layers in Algorithm 2, not visiting every exponentially growing linear region [27, 18]. We note that this formal explanation through tropical geometry is unprecedented in the previous work [4].
@@ 3.2 Edge subdivision for tropical geometry
for all $i$ and $j$ where $\phi(i,j)\leq\phi(i^{*},j^{*})$, and $\phi(i,j)$ is an ascending indexing function from lower layers, to vectorize a signed matrix, handling arbitrary hidden sizes of networks.
@@ 4.1 Piecewise trilinear networks
Let a unit cube be in the $D$-dimensional space, where its corners are represented by ${\bm{H}}\in\mathbb{R}^{2^{D}\times F}$. Given weights ${\bm{w}}\in[0,1]^{D}$, the $\psi$ is defined as:
@@
where the interpolating weight $\omega(i,{\bm{w}})\in\mathbb{R}$ is defined using a left-aligned and zero-trailing binarization function $\textrm{B}\in\{0,1\}^{D}$ (ref., Section B.3) as follows:
@@
which is the volume of the opposite subsection of the hypercube divided by the weight point ${\bm{w}}$. For a general hypercube, scaling the weight for a unit hypercube gets the same result. Trilinear interpolation extends linear and bilinear interpolations to three dimensions, where $D=3$.
@@
Let a positional encoding module be $\tau:\mathbb{R}^{D}\rightarrow\mathbb{R}^{F}$ using the trilinear interpolation of spatially nearby learnable vectors on a three-dimensional grid, e.g., HashGrid [12] or TensoRF [13]. Usually, they transform the input coordinate ${\bm{x}}$ to ${\bm{w}}$ with a multi-resolution scale and find a relative position within a grid cell. Then, we define trilinear neural networks $\tilde{\nu}$, by prepending $\tau$ to the neural networks $\nu$ from Definition 3.2. Here, the input and output dimensions of $\nu^{(L)}$ are $F$ and $1$ (an SDF distance), respectively.
@@
Note that $\tau$ is grid-wise trilinear. Using Lemma 4.2, $\tilde{\nu}$ is trilinear within the composite intersections of linear regions of $\nu$ and trilinear regions of $\tau$ if available. (We will discuss how to access the cubic corner features in the case that trilinear regions are no longer cubic in Section 5.) Then, we regard $\tilde{\nu}$ as piecewise trilinear.
@@ 4.2 Curved edge subdivision in a trilinear space
Notice that it does not decrease the number of variables to find a line solution since the trilinear interpolation $\tau$ forms hypersurfaces making complex cases. Yet, we theoretically demonstrate that the eikonal constraints on $\tilde{\nu}$ render them hyperplanes with linear solutions.
@@
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
Let $\nu_{0}({\bm{x}})$ and $\nu_{1}({\bm{x}})$ be two affine transformations defined by ${\bm{W}}_{i}^{\intercal}{\bm{x}}+{\bm{b}}_{i}$, $i\in\{0,1\}$. A hypersurface $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ passing two points $\nu_{0}\big{(}\tau({\bm{x}}_{0})\big{)}=\nu_{0}\big{(}\tau({\bm{x}}_{7})\big{)}=0$ while $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}\neq 0$ satisfies the eikonal constraint $\|\nabla\nu_{1}\big{(}\tau({\bm{x}})\big{)}\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ is a plane.
@@
Since we aim for a practical polyhedral complex derivation that piecewise (still) trilinear networks represent, we replace one of the hypersurfaces forming the curved edge with a diagonal plane in a piecewise trilinear region. We opt for this choice not only due to its mathematical simplicity but also the characteristics of trilinear hypersurfaces as illustrated in Figure 4, Appendix G. Thus, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error from this approximation. This error is tolerated by the hyperparameter ϵ=italic-$\epsilon=$1e-4 as specified in Definition 3.4 and Section 6. More discussions on this matter can be found in Section B.1.
@@
while
@@ 4.3 Skeletonization, faces, and normals
Skeletonization selects the vertices such that: ${\mathcal{V}}^{\star}=\{{\bm{x}}|\tilde{\nu}({\bm{x}})\leq\epsilon,{\bm{x}}\in{\mathcal{V}}\}$, and the edges ${\mathcal{E}}^{\star}$ such that their two vertices are among ${\mathcal{V}}^{\star}$.
Edges lying on the same plane and a shared region form a face, and this alignment can be determined by evaluating the sign-vectors (ref. Definition 3.4 and Section 5). Note that faces are not inherently triangular; however, if needed, they can be triangulated via triangularization for further analysis. The order of vertices conventionally describes its normal. Let ${\bm{x}}_{0}$, ${\bm{x}}_{1}$, and ${\bm{x}}_{2}$ be the vertices forming a triangular face. The normal is defined using cross-product as ${\bm{n}}=({\bm{x}}_{1}-{\bm{x}}_{0})\times({\bm{x}}_{2}-{\bm{x}}_{0})$ where the normal is orthogonal to both vectors, with a direction given by the right-hand rule. If the viewing direction aligns opposite to the normal, the order of vertices follows a counter-clockwise arrangement. We determine the vertex order based on the normal derived from the Jacobian of the mean of vertices. Using cosine similarity to compute relative angles $\theta\in[-\pi,\pi]$ and cross product for getting direction, we identify the counter-clockwise arrangement of vertices, as outlined in Algorithm 3 in Appendix.

### v3

# Jin-Hwa Kim NAVER AI Lab & SNU AIIS Republic of Korea

@@ Abstract
Recent advancements in visualizing deep neural networks provide insights into their structures and mesh extraction from Continuous Piecewise Affine (CPWA) functions. Meanwhile, developments in neural surface representation learning incorporate non-linear positional encoding, addressing issues like spectral bias; however, this poses challenges in applying mesh extraction techniques based on CPWA functions. Focusing on trilinear interpolating methods as positional encoding, we present theoretical insights and an analytical mesh extraction, showing the transformation of hypersurfaces to flat planes within the trilinear region under the eikonal constraint. Moreover, we introduce a method for approximating intersecting points among three hypersurfaces contributing to broader applications. We empirically validate correctness and parsimony through chamfer distance and efficiency, and angular distance, while examining the correlation between the eikonal loss and the planarity of the hypersurfaces. The code is available at https://github.com/naver-ai/tropical-nerf.pytorch.
@@ 1 Introduction
Recent advancements in visualizing deep neural networks [1, 2, 3, 4] significantly contribute to understanding their intricate structures. This progress provides valuable insights into the expressivity, robustness, training methodologies, and distinctive geometry of neural networks. By leveraging the inherent piecewise linearity in the Continuous Piecewise Affine (CPWA) functions, e.g., ReLU neural networks, each region of the input space is represented as a convex polyhedron. The assembly of these sets constructs a polyhedral complex that delineates the decision boundaries of neural networks [5].
@@
    We provide a theoretical analysis of the piecewise trilinear networks under the eikonal constraint revealing that, within a trilinear region, a hypersurface transforms into a plane.
@@ 2 Related work
A conventional black-box approach is to assess a given function at sampled points to derive the complex of the function, commonly through marching cubes [9] or marching tetrahedra [16, 10]. Despite the inclusion of mesh optimization techniques [17] in modern mesh packages, this exhaustive method often generates redundant complexities due to discretization, increasing the computational load by using unnecessary mesh elements (e.g., fragmented planars). Aiming for exact extraction, initial efforts focused on identifying linear regions [18] and employing Analytical Marching [2] to exactly reconstruct the zero-level isosurface. Contemporary approaches involve region subdivision, wherein the regions of the complex are progressively subdivided from neuron to neuron and layer to layer, allowing for the efficient calculation of the exponentially increasing linear regions [19, 20, 21, 4]. Nevertheless, many previous studies have explored CPWA functions consisting of fully connected layers and ReLU activation, primarily owing to the mathematical simplicity of hyperplanes defining linear regions. However, this restricts their applications in emerging domains of deep implicit surface learning that diverge from CPWA functions.
@@
The spectral bias of a multi-layer perceptron (MLP) hinders effectively learning high frequencies, both theoretically and empirically [14]. To address this limitation, Fourier feature mapping, employing trigonometric projections similar to those used in the Transformer architecture [22], proves successful in representing complex 3D objects and scenes. Furthermore, to address its slow convergence and enhance rendering quality, novel positional encoding methods using trilinear interpolations, e.g., HashGrid [12] or TensoRF [13], are introduced. Both techniques generate positional features through trilinear interpolation among the eight nearest corner features derived from pre-defined 3D grids. These features are obtained by hashing from multi-resolution hash tables or factorized representations of a feature space, respectively. However, these approaches cause the function represented by neural networks to deviate from CPWA functions, as trilinear interpolation is not an affine transformation. Consequently, the complex regions are divided by intricate hypersurfaces.
@@
An eikonal  [^1] The eikonal is derived from the Greek, meaning image or icon. equation is a non-linear first-order partial differential equation governing the propagation of wavefronts. Notably, this finds application in the regularization of a signed distance function (SDF) in the context of neural surface modeling [23, 6, 8]. If $\Omega$ represents a subset of the space ${\mathcal{X}}$ equipped with the metric $d$, the SDF $f({\bm{x}})$ is defined as follows:
@@
where $\partial\Omega$ denotes the boundary of $\Omega$. The metric $d$ is defined (with a slight notational abuse) as:
@@ 3 Preliminaries
We present an overview of tropical geometry (Section 3.1) to provide a formal definition of the tropical hypersurface and the tropical algebra of neural networks. Then, we discuss the edge subdivision algorithm [4] for extracting the tropical hypersurface (Section 3.2). We supplement Section A.1 for less familiar readers. Later, we extend this idea to include considerations for trilinear interpolation in Section 4. You may choose to skip these sections if you are already acquainted with these concepts.
@@ 3.1 Tropical geometry
Tropical  [^2] The tropical is named in honor of Hungarian-born Brazilian computer scientist Imre Simon [24]. geometry [25, 26] describes polynomials and their geometric characteristics. As a skeletonized version of algebraic geometry, tropical geometry represents piecewise linear meshes using tropical semiring and tropical polynomial (ref. Section A.1). The set of points where a tropical polynomial $f$ is non-differentiable, due to tropical sum $\oplus:=\max(\cdot)$, is called tropical hypersurface:
@@
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c_{i}\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$ and tropical product :=$\odot$:=+⊙ := +.
@@
Assuming input and output dimensions are consistent, $L$-layer neural networks with ReLU activation are a composition of functions defined as:
@@
Although the neural networks $\nu$ is generally non-convex, it is known that the difference of two tropical signomials (allowing a real-valued $\alpha$ in a tropical polynomial) can represent $\nu$ [27] (ref. Section A.3). Given this observation, we proceed to examine the following proposition:
Proposition 3.3 (Decision boundary of neural networks).
The decision boundary for $L$-layer neural networks with ReLU activation $\nu^{(L)}({\bm{x}})=0$ is a subset of or equals with the region where two tropical monomials, or two arguments of $\max$, equal (Definition 3.1), as follows:
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)~{}\textnormal{if}~{}i\neq L~{}\textnormal{else}~{}(1)\big{\}}$
where the subscript $j$ denotes the $j$-th neuron, assuming the hidden size of neural networks is $H$, while the final $L$-th layer has a single output for rather discussions as an SDF (Definition 4.3). Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
The piecewise hyperplanes corresponding to $\nu_{j}^{(i)}$ are the candidates shaping the decision boundary. It enables for-loop iterations over the neurons across layers, as in Algorithm 2, not visiting every exponentially growing linear region [27, 18]. The equality in Proposition 3.3 generally does not hold, but we can filter them by checking if $\nu_{1}^{(L)}({\bm{x}})=0$. For more details, please refer to Berzins [4]. We note that our formal explanation through tropical geometry is unprecedented in the previous work [4].
@@ 3.2 Edge subdivision for tropical geometry
for all $i$ and $j$ where $\phi(i,j)\leq\phi(i^{*},j^{*})$, and $\phi(i,j)\in\{0\}\cup{\mathbb{N}}$ is an ascending indexing function from lower layers, handling arbitrary hidden sizes of networks to vectorize a signed matrix.
@@ 4.1 Piecewise trilinear networks
Let a unit cube be in the $D$-dimensional space, where its corners are represented by ${\bm{H}}\in\mathbb{R}^{2^{D}\times F}$ where $F$ is a corner-feature dimension. Given weights ${\bm{w}}\in[0,1]^{D}$, the trilinear interpolation function $\psi$ is defined as:
@@
where the interpolating weight $\omega(i,{\bm{w}})\in\mathbb{R}$ is defined using a left-aligned and zero-trailing binarization function $\textrm{B}\in\{0,1\}^{D}$ (ref. Section B.1) as follows:
@@
which is the volume of the opposite subsection of the hypercube divided by the weight point ${\bm{w}}$. For a general hypercube, scaling the weight for a unit hypercube gets the same result. Going forward, we assume $D$=3 for trilinear interpolation without explicitly stating it otherwise.
@@
Let a positional encoding module be $\tau:\mathbb{R}^{D}\rightarrow\mathbb{R}^{F}$ using the trilinear interpolation of spatially nearby learnable vectors on a three-dimensional grid, e.g., HashGrid [12] or TensoRF [13], where $\tau$ is an instance of $\psi$ in Definition 4.1. Usually, they transform the input coordinate ${\bm{x}}$ to ${\bm{w}}$ as a relative position inside a unit grid in a corresponding resolution. ${\bm{H}}$ is a learnable parameter specified in the corresponding method. Then, we define trilinear neural networks $\tilde{\nu}$, by prepending $\tau$ to the neural networks $\nu$ from Definition 3.2. Here, the input and output dimensions of $\nu^{(L)}$ are $F$ and $1$ (an SDF distance), respectively.
@@
Note that $\tau$ is gridwise trilinear. Using Lemma 4.2, $\tilde{\nu}$ is trilinear within the composite intersections of linear regions of $\nu$ and trilinear regions of $\tau$. (We discuss how to access virtual cubic corner features where trilinear regions are not cubic in Section 5.) So, we regard $\tilde{\nu}$ as piecewise trilinear.
@@ 4.2 Curved edge subdivision in a trilinear space
Notice that it does not decrease the number of variables to find a line solution since the trilinear interpolation $\tau$ still forms hypersurfaces making complex cases. Yet, we theoretically demonstrate that the eikonal constraints on $\tilde{\nu}$ render them hyperplanes with linear solutions.
@@
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}^{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
Let $\nu_{0}({\bm{x}})$ and $\nu_{1}({\bm{x}})$ be two affine transformations defined by ${\bm{W}}_{i}^{\intercal}{\bm{x}}+{\bm{b}}_{i}$, $i\in\{0,1\}$. A hypersurface $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ passing two points $\nu_{0}\big{(}\tau({\bm{x}}_{0})\big{)}=\nu_{0}\big{(}\tau({\bm{x}}_{7})\big{)}=0$ while $\nu_{0}\big{(}\tau({\bm{x}}_{1\dots 6})\big{)}\neq 0$ satisfies the eikonal constraint $\|\nabla\nu_{1}\big{(}\tau({\bm{x}})\big{)}\|_{2}^{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\nu_{0}\big{(}\tau({\bm{x}})\big{)}=0$ is a plane.
@@
Since we aim for a practical polyhedral complex derivation that piecewise trilinear networks represent, we replace one of the hypersurfaces forming the curved edge with a diagonal plane in a piecewise trilinear region. We choose this option not only for its mathematical simplicity but also due to the characteristics of trilinear hypersurfaces, as illustrated in Figure 9, Appendix G. Thus, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error from this approximation. This error is tolerated by the hyperparameter ϵ=italic-$\epsilon=$1e-4 as specified in Definition 3.4 and Section 6. More discussions on this matter can be found in Section C.1.
@@
${\bm{X}}^{\intercal}\big{(}{\bm{P}}_{\alpha}{\bm{Q}}_{\beta}^{\intercal}-{\bm{P}}_{\beta}{\bm{Q}}_{\alpha}^{\intercal}\big{)}{\bm{X}}=0,~{}~{}~{}~{}~{}\text{while}~{}~{}~{}~{}y$ $=\frac{{\bm{X}}^{\intercal}{\bm{P}}_{\alpha}}{{\bm{X}}^{\intercal}({\bm{P}}_{\alpha}-{\bm{P}}_{\beta})}\hskip 10.00002pt({\bm{P}}_{\alpha}\neq{\bm{P}}_{\beta}).$ (8)
@@ 4.3 Skeletonization, faces, and normals
Skeletonization selects the vertices such that: ${\mathcal{V}}^{\star}=\{{\bm{x}}~{}|~{}|\tilde{\nu}({\bm{x}})|\leq\epsilon,{\bm{x}}\in{\mathcal{V}}\}$, and the edges ${\mathcal{E}}^{\star}$ such that their two vertices are among ${\mathcal{V}}^{\star}$. Recall that because a decision boundary is a subset of the tropical hypersurface (Definitions 3.1 and 3.3), this step ensures accurate mesh extraction.
Faces are formed by edges lying on the same plane and sharing a common region. This is identified by evaluating the sign-vectors (ref. Definition 3.4 and Section 5). Note that faces are not inherently triangular; however, if needed, they can be triangulated via triangularization for further analysis.
Normals are conventionally described by the order of the vertices as each face has ambiguity with two opposite directions. Let ${\bm{x}}_{0,1,2}$ be the vertices forming a triangular face. The normal is defined using cross-product as ${\bm{n}}=({\bm{x}}_{1}-{\bm{x}}_{0})\times({\bm{x}}_{2}-{\bm{x}}_{0})$ where the normal is orthogonal to both vectors, with a direction given by the right-hand rule. Please refer to Section B.2 for the details.
Figure 2: Trilinear regions in the xy-plane at $z=0.04$, identified by the sign-vectors (Definition 3.4), are represented with random colors. (a) Grids described in Section 5 and Algorithm 4. (b) The neurons of the first layer representing folded hypersurfaces (blue arrow). (c) All neurons representing every nonlinear boundary. (d) Select all zero-set vertices and edges. (e) Skeletonized as in Section 4.3.